OpenStudy (anonymous):

A large soda bottle cap has a radius of 1+2√2 centimeters. What is the area of the bottle cap? (Recall that the area of a circle is πr^2 where r is the radius).

OpenStudy (anonymous):

@johnweeden

OpenStudy (amistre64):

they give you a formula, and even tell you the value of r ....

OpenStudy (amistre64):

what part is giving you trouble?

OpenStudy (anonymous):

cenimeters

OpenStudy (amistre64):

centimeters is just the unit measure of length, and gets attached to the r. cm*cm = cm^2

OpenStudy (anonymous):

(9+8√2)π square cenimeters

OpenStudy (amistre64):

(1+2sqrt(2))^2 pi squre centimeters 1 + 8 + 4sqrt2 if im seeing the squaring correctly

OpenStudy (amistre64):

yeah 4sqrt2, not 8sqrt2

OpenStudy (anonymous):

its (9+4√2)π square cenimeters thnk

OpenStudy (amistre64):

yeah, that should be it :)

OpenStudy (anonymous):

6\[6\sqrt{2}-15\sqrt{12}\] is what i got